OpenStudy (anonymous):

Circle S has an equation of (x – 16)2 + (y + 9)2 = 4. What is the center and radius of circle S?

OpenStudy (anonymous):

x – 16)^{2} + (y + 9)^{2} = 4.

OpenStudy (anonymous):

U know you can just read this from the equation, right?

OpenStudy (anonymous):

(x - h)2 + (y - k)2 = r2

OpenStudy (anonymous):

i know that,but i was trying to fix the powers,thanks.

OpenStudy (amistre64):

the powers are superfluous ....

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

\[\huge (x-\underbrace{h}_{Center})^2+(y-\underbrace{k}_{Center})^2=\underbrace{\boxed{r}^2}_{Radius }\]

OpenStudy (amistre64):

im sure its in the dictionary

OpenStudy (anonymous):

okay im not talking to you right now,thanks.

OpenStudy (amistre64):

lol, the powers are not important to determing the center from the equation is all

OpenStudy (anonymous):

okay i obviously put the question on here for a reason,because i didn't know what it meant.thus,i pasted it on here for someone other than myself to figure it out.

OpenStudy (anonymous):

So did Kreshnik's answer help you?

OpenStudy (anonymous):

yes,it did.